OpenStudy (anonymous):

Choose the correct simplification of the expression (a2)3. a5 a8 a a6

OpenStudy (anonymous):

The fourth one

OpenStudy (anonymous):

Thanksssss! :D

jimthompson5910 (jim_thompson5910):

please don't just give out answers give out steps to get there

OpenStudy (anonymous):

the rule here is to multiply the exponents

jimthompson5910 (jim_thompson5910):

(a^2)^3 is really saying (a^2)*(a^2)*(a^2) = a^(2+2+2) = a^6 or you can just multiply the exponents to get a^(2*3) = a^6

OpenStudy (anonymous):

Ohhh.

OpenStudy (anonymous):